OpenStudy (anonymous):

i need help with 3 others like this including this one: 6. Use the polynomial to answer the question. 6b3 + 9b2 - 12b Which represents the polynomial in factored form? (Points : 3) 3b3(2b2 + 3b - 4) 3b(2b2 + 9b2 - 12b) 3b2(3b2 + 6b - 9) 3b(2b2 + 3b - 4)

OpenStudy (jusaquikie):

for factoring you are just pulling out the most that you can get from each term, start with the variables, they are easiest.\[b(6b^2+9b+12)\] there is no more like variables in all 3 terms. now move to the numbers, what will 6, 9, and 12 all divide by?

OpenStudy (anonymous):

3

OpenStudy (jusaquikie):

should be a -12 up there \[b(6b+9b-12)\]

OpenStudy (anonymous):

6,912 are divisible by 3

OpenStudy (anonymous):

*6,9,12

OpenStudy (jusaquikie):

so pull the 3 out and divide each number by 3 and you have your answer

OpenStudy (anonymous):

is it that easy? they make it so hard in the actual lesson 6/3 = 2 9/3 = 3 -12/3 = -4

OpenStudy (anonymous):

so the answer is a?

OpenStudy (jusaquikie):

how many b's did i pull out?

OpenStudy (anonymous):

3

OpenStudy (jusaquikie):

lol ok to what power did i pull b out

OpenStudy (anonymous):

^3

OpenStudy (anonymous):

or is that just the degree?

OpenStudy (jusaquikie):

b^3 = b*b*b

OpenStudy (jusaquikie):

the last term only had 1 b so how could i pull out b*b*b?

OpenStudy (anonymous):

no because the degree is the highest variable

OpenStudy (anonymous):

are you saying you have b^2 * b?

OpenStudy (anonymous):

because u pulled them out of the ()

OpenStudy (jusaquikie):

factoring is just rearranging terms to find out other ways to write them just like 20 = 1*20=10*2=4*5 and B^3=B^2*b=b*b*b

OpenStudy (anonymous):

i see

OpenStudy (jusaquikie):

so one way to check your answer is to multiply each term by what's outside the parentheses and see if if gives you back the original function.

OpenStudy (anonymous):

i see so: the answer is d because: 3b x 2b^2 = 6b^3 3b x 3b = 9b^2 and 3b x -4 = -12b

OpenStudy (jusaquikie):

yes and just like you multiplied to check your answer when your trying to factor without a multiple choice you just want to find the biggest thing that all 3 are divisible by

OpenStudy (anonymous):

7. Use the polynomial to answer the question. 16x2 - 25 Which represents the polynomial in factored form? (Points : 3) (4x + 5)(4x - 5) (4x 5)(4x- 5) (8x 5)(8x -5) (8x + 5)(8x - 5)

OpenStudy (jusaquikie):

but it's easiest to do it in steps, look for likenesses in all 3 terms, the lowest order is usually the easiest because if it has only 1 variable then that's all you can factor out. if it has none then you can only look for the greatest common factor

OpenStudy (anonymous):

i went with a (used the foil method) 4x * 4x = 6x^2 4x x -5 = -25x 5 x 4x = 25x 5 x - 5 = -25 so i would have : 6x^2 -25x +25x -25 6x^2 - 25

OpenStudy (jusaquikie):

ok once you learn some a few tricks this form is the easiest and you will see why, this one is missing the middle term and it's last term is negative. take the square of each term, then write that twice one with a (ax+b) and one with a (ax-b) that way when they multiply back out the abx and the -abx cancel out and disapear

OpenStudy (anonymous):

like -25 +25?

OpenStudy (jusaquikie):

no when you are multiplying polynomials you have to multiply all out so it would be like this

OpenStudy (jusaquikie):

i'm going to do the first answer (4x+5)(4x-5) which happens to be the answer lol so first (4x*4x) then (4x*-5) then (5*4x) then (5*-5) so we have (16x²)+(-20x)+(20x)+(-25) add those all up

OpenStudy (anonymous):

you go to love the foil method lol

OpenStudy (anonymous):

so then i would be able to do the last one to right? it is: Use the polynomial to answer the question. x2 - 8x + 15 Which represents the polynomial in factored form? (Points : 3) (x + 3)(x 5) (x + 3)(x + 5) (x - 3)(x + 5) (x - 3)(x - 5)

OpenStudy (anonymous):

1 would go into all

OpenStudy (anonymous):

2x 4 = 8 and 3 x 5 = 15 thats why i thought it was suposed to look like this (x^2 -(8x) +15) or (x^2 -8x) and (8x +15)

OpenStudy (jusaquikie):

these last 2 are diffrent from the first one, with these you are liiking for something in the form (ax+b)(cx+d) or (ax+b)²

OpenStudy (anonymous):

so that means that foil would not work this time how would i go about solving the problem itself?

OpenStudy (jusaquikie):

foil will take one of the answer set to the question

OpenStudy (anonymous):

because (-) is first and (=) is second in the original problem: (x - 3)(x + 5) (x * x) = x^2 ( x * 5) = 5x (-3 * x) = - 3x (-3 * 5 = -15 x^2 + 5x - 3x - 15? x^2 + 2 - 15

OpenStudy (jusaquikie):

here is how i factor polynomials, we are looking for our answer in a form like this (x±a)(x±b)

OpenStudy (anonymous):

because that was not the answer i would say "b" would be it if not "a":

OpenStudy (jusaquikie):

you con't subtract the x but the rest is right it would be x²+2x-15

OpenStudy (anonymous):

but that is not the original problem because of that i would say the one with both minus would be it

OpenStudy (jusaquikie):

ok we want a positive 15 but we want the two x terms to add to -8 what would we have to multiple to get something that would add to -8x?

OpenStudy (anonymous):

because 2 negative when multiplied make a positive: (x - 3)(x-5) x^2 - 5x- 3x+15 x^2 - 8 + 15

OpenStudy (jusaquikie):

ok so let's get some intuition looking at the last term (15) it's positive and it's only effected by the b part in (x+b) so for it to be positive it has to have 2 negatives or 2 positives multiplied together

OpenStudy (jusaquikie):

i think you have it now

OpenStudy (anonymous):

and i cant be 2 positives because then the 8 would not fit

OpenStudy (jusaquikie):

correct

OpenStudy (anonymous):

thank you :)

OpenStudy (jusaquikie):

np